OpenStudy (anonymous):

Solve using the additional principal don't forget to perform a check x+2/13=7/13 x=?

hero (hero):

x + 2 = 7 x = 7 - 2 x = 5 (x+2)/13 = 7/13 (5+2)/13 = 7/13 7/13 = 7/13

hero (hero):

Any questions?

OpenStudy (anonymous):

6x+13x=152 x=?

hero (hero):

Candace, what grade are you in?

OpenStudy (anonymous):

I'm taking college algebra